OpenStudy (anonymous):

simplify (((1/(5+h))-(1/5))/h

OpenStudy (anonymous):

no its (((1/(5+h))-(1/5)) and that is all over h so the numerator if the fraction is (((1/(5+h))-(1/5)) and the denominator is h

OpenStudy (anonymous):

(1/(5+h))-(1/5) = (5 - (5+ h) / 5(5+ h) = -h / 5(5+ h) now divide by h: = -1 [(5(5+h)]

OpenStudy (curry):

-25-h

OpenStudy (anonymous):

sorry - i left out the divide: '/' correct is -1 / [(5(5+h)]